# The Effects of Multi-Task Learning on ReLU Neural Network Functions

### v2

@@ Abstract
This paper studies the properties of solutions to multi-task shallow ReLU neural network learning problems, wherein the network is trained to fit a dataset with minimal sum of squared weights. Remarkably, the solutions learned for each individual task resemble those obtained by solving a kernel method, revealing a novel connection between neural networks and kernel methods. It is known that single-task neural network training problems are equivalent to minimum norm interpolation problem in a non-Hilbertian Banach space, and that the solutions of such problems are generally non-unique. In contrast, we prove that the solutions to univariate-input, multi-task neural network interpolation problems are almost always unique, and coincide with the solution to a minimum-norm interpolation problem in a Sobolev (Reproducing Kernel) Hilbert Space. We also demonstrate a similar phenomenon in the multivariate-input case; specifically, we show that neural network learning problems with large numbers of diverse tasks are approximately equivalent to an $\ell^{2}$ (Hilbert space) minimization problem over a fixed kernel determined by the optimal neurons.
@@ 1 Introduction
This paper investigates the nature of functions learned by training neural networks on multiple “tasks” (i.e., multiple sets of labels on the same set of data points) and contrasts these solutions with those obtained by learning separate networks for each individual task. We show that the solutions of multi-task neural network learning problems can differ dramatically from the solutions to single-task problems, even if the tasks are diverse and unrelated to one another. Unlike standard intuitions Caruana (1997) and existing theory Ben-David and Schuller (2003); Maurer et al. (2016) regarding the effects and benefits of multi-task learning, our results do not rely on similarity between tasks.
@@
Neural networks are trained to fit data using gradient descent methods and often include a form of regularization called weight decay, which penalizes the $\ell^{2}$ norm of the network weights. We consider weight decay applied only to the input and output weights of the neurons—no regularization is applied to the biases or residual connection. This is a common setting studied frequently in past work Savarese et al. (2019); Ongie et al. (2019); Parhi and Nowak (2021). Intuitively, only the input and output weights—not the biases or residual connection—affect the “regularity” of the neural network function as measured by its second (distributional) derivative, which is why it makes sense to regularize only these parameters. Given a set of training data points $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$ and a fixed width $K\geq N^{2}$, we consider the weight decay interpolation problem:  [^1] By an argument similar to the proof of Theorem 5 of Shenouda et al. (2024), the optimal objective value of (3) (hence of (2)) for any given dataset is the same for any $K\geq N^{2}$; this is the scenario that we consider in this work.
@@
The optimizations in (2) and (3) are non-convex and in general, they may have multiple global minimizers. As an example, consider the single-task, univariate dataset in Figure 1. For this dataset, (3) has infinitely many global solutions Savarese et al. (2019); Ergen and Pilanci (2021); Debarre et al. (2022); Hanin (2021). Two of the global minimizers are shown in Figure 1. In some scenarios, the solution on the right may be preferable to the one on the left, since the interpolation function stays closer to the training data points, and hence is more adversarially robust by most definitions Carlini et al. (2019). Moreover, recent theoretical work shows that this solution has other favorable generalization and robustness properties Joshi et al. (2024). Current training methods, however, might produce any one of the infinite number of solutions, depending on the random initialization of neural network weights as well as other possible sources of randomness in the training process. It is impossible to control this using existing training algorithms, which might explain many problems associated with current neural networks such as their sensitivity to adversarial attacks. In contrast, as we show in this paper, if we train a neural network to interpolate the data in Figure 1 along with additional interpolation tasks with different labels, then the solution is almost always unique and given by the (potentially preferable) interpolation depicted on the right. This demonstrates that multi-task learning can have a profound effect on the nature of neural network solutions.
@@
  2. Multi-task Training $\equiv$ Kernel Method (almost always).
    In the case that the solution to the univariate weight decay problem is unique, it is given by the connect-the-dots interpolant of the training data points: i.e., the optimal solution is a linear spline which performs straight-line interpolation between consecutive data points in all tasks. On the support of the data, this solution agrees with the solution to minimum-norm interpolation in the first-order Sobolev space $H^{1}$, a reproducing kernel Hilbert space (RKHS) which contains all functions with first derivatives in $L^{2}$ De Boor and Lynch (1966). In contrast, solutions to the single-task learning problem are non-unique in general and are given by minimum-norm interpolating functions in the non-Hilbertian Banach space $\operatorname{BV}^{2}$ Parhi and Nowak (2021), which contains all functions with second distributional derivatives [^2] Technically, $\operatorname{BV}^{2}$ contains all functions with second distributional derivatives in ${\mathcal{M}}$, the space of Radon measures with finite total variation. ${\mathcal{M}}$ can be viewed as a “generalization” of $L^{1}$ (see Ch. 7.3, p.223 in Folland (1999)). in $L^{1}$. This shows that the solution to each individual task in a multi-task solution is equivalent to that of a kernel method, whereas single-task solutions generally are not.
  3. Insights on Multivariate Multi-Task Problems.
    We provide empirical evidence and mathematical analysis which indicate that similar conclusions hold in multivariate settings. Specifically, with a large number of diverse tasks, the solutions for each individual task in multi-task learning are approximately minimum-norm solutions in a particular RKHS space determined by the optimal neurons. In contrast, learning each task in isolation results in solutions that are minimum-norm with respect to a non-Hilbertian Banach norm over the optimal neurons.
@@ 2 Related Works
For single-output ReLU neural networks, Savarese et al. (2019); Ongie et al. (2019) related weight decay regularization on the parameters of the model to regularizing a particular semi-norm on the neural network function. Ongie et al. (2019) showed that this semi-norm is not an RKHS semi-norm, highlighting a fundamental difference between neural networks and kernel methods. Parhi and Nowak (2021, 2022); Bartolucci et al. (2023); Unser (2021) studied the function spaces associated with this semi-norm, and developed representer theorems showing that optimal solutions to the minimum-norm data fitting problem over these spaces are realized by finite-width ReLU networks. Consequently, finite-width ReLU networks trained with weight decay are optimal solutions to the regularized data-fitting problem posed over these spaces. Function spaces and representer theorems for multi-output and deep neural networks were later developed in Korolev (2022); Parhi and Nowak (2022); Shenouda et al. (2024).
Characterizations of ReLU neural network solutions:
Hanin (2021); Stewart et al. (2023) characterized the neural network solutions to (2) in the univariate input/output setting. Boursier and Flammarion (2023) showed that in the univariate input/output case, when weight decay is modified to include the biases of each neuron the solution is unique. Moreover, under certain assumptions it is often the sparsest interpolator (in terms of the number of neurons). Our work differs from these as we study the multi-task setting, showing that the solutions are almost always unique and that they are the connect-the-dots solution, which is generally not the sparsest and corresponds to minimum norm interpolation over a particular RKHS space. While characterizing solutions to (2) in the multivariate setting is more challenging, there exist some results for very particular datasets Ergen and Pilanci (2021); Ardeshir et al. (2023); zeno2024minimum.
@@
The advantages of multi-task learning has been extensively studied in the machine learning literature Obozinski et al. (2006, 2010); Argyriou et al. (2006, 2008). Theoretical properties of multi-task neural networks in particular, have also been studied in Lindsey and Lippl (2023); Collins et al. (2024); Shenouda et al. (2024). The underlying intuition in these past works has been that learning multiple related tasks simultaneously can help select or learn the most useful features for all tasks. Our work differs from this traditional paradigm as we consider multi-task neural networks trained on very general tasks which may be diverse and unrelated.
@@ 3 Univariate Multi-Task Neural Network Solutions
The “representational cost” of a function $f$ that can be represented as a neural network (1) with unit-norm input weights, i.e. $\|{\bm{w}}_{k}\|_{2}=1$ for all $k=1,\dots,K$, is defined to be
$R(f):=\inf_{\bm{\theta}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}\ \quad\text{s.t.}\quad f=f_{{\bm{\theta}}}$ (4)
@@
denotes the slope of the straight line between $(x_{i},y_{it})$ and $(x_{i+1},y_{i+1,t})$. The connect-the-dots interpolant of the dataset ${\mathcal{D}}_{t}$ is the function $f_{{\mathcal{D}}_{t}}$ which connects the consecutive points in dataset ${\mathcal{D}}_{t}$ with straight lines (see Figure 2). Its slopes on $(-\infty,x_{2}]$ and $[x_{N-1},\infty)$ are $s_{1t}$ and $s_{N-1,t}$, respectively.
@@
This result raises an interesting new connection between data fitting with ReLU neural networks and traditional kernel-based learning methods, because connect-the-dots interpolation is also the minimum-norm solution to a data-interpolation problem in the first-order Sobolev space $H^{1}([x_{1},x_{N}])$, itself an RKHS whose norm penalizes the $L^{2}$ norm of the function’s derivative. In particular, $f_{{\mathcal{D}}_{t}}$ agrees on $[x_{1},x_{N}]$ with the function $f(x)=\sum_{j=1}^{N}\alpha_{j}k(x,x_{j})$ whose coefficients $\alpha_{j}$ solve the kernel optimization problem
@@
Previous works Shenouda et al. (2024) and Lindsey and Lippl (2023) showed that multi-task training encourages neuron sharing, where all task are encouraged to utilize the same set of neurons or representations. Our result shows that univariate multi-task training is an extreme example of this phenomenon, since $f_{{\mathcal{D}}}$ can be represented using only $N-1$ neurons, all of which contribute to all of the network outputs. Therefore, in the scenario we study here, neuron sharing almost always occurs even if the tasks are unrelated.
@@ 3.1 Characterization and Uniqueness
The full proof of Theorem 3.1 appears in Appendix 6.1. We outline the main ideas here. Our proof relies on the fact that any $\mathbb{R}\to\mathbb{R}^{T}$ ReLU neural network of the form (1) which solves (3) represents $T$ continuous piecewise linear (CPWL) functions, where the change in slope of the $t^{\textrm{th}}$ function at the $k^{\text{th}}$ knot is equivalent the $t^{\textrm{th}}$ entry of the $k^{\text{th}}$ output weight vector (see Appendix 6.1 for further detail). This fact allows us to draw a one-to-one correspondence between each knot in the function and each neuron in the neural network. The proof relies primarily on the following lemma
@@
[…]
@@
Figure 3: The $t^{\text{th}}$ neural network output $f_{{\bm{\theta}}t}$ around the knot at $x$, where $\tau=\frac{x-x_{i}}{x_{i+1}-x_{i}}$. Each line segment in the figure is labeled with its slope. For any particular output $t$, it may be the case that $f_{{\bm{\theta}}t}$ does not have a knot at $x$ (in which case $\delta_{t}=0$); does not have a knot at $x_{i}$ (in which case $a_{t}=b_{t}+\delta_{t}$); and/or does not have a knot at $x_{i+1}$ (in which case $b_{t}-\frac{\tau}{1-\tau}\delta_{t}=c_{t}$).
@@
The representational cost (4) is separable across neurons. The contribution of these neurons to $R(f)$ is:
@@
where the inequality follows from the triangle inequality. This shows that taking $\delta_{t}=0$ for all outputs, which corresponds to connecting $x_{1}$ and $x_{2}$ with a straight line in all outputs, will never increase the representational cost of $f$. The triangle inequality used in (3.1) holds with equality for some $\bm{\delta}\neq\bm{0}$ if and only if ${\bm{a}}-{\bm{b}}$, ${\bm{b}}-\bm{c}$, and $\bm{\delta}$ are aligned with $\|\bm{\delta}\|_{2}\leq\min\{\|{\bm{a}}-{\bm{b}}\|_{2},\frac{1-\tau}{\tau}\|{\bm{b}}-\bm{c}\|_{2}\}$. ∎
@@
Lemma 3.2 states that removing neurons which are located away from the data points and replacing them with a straight line will never increase the representational cost of the network, and it will strictly decrease the representational cost unless ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned. This result implies that the connect-the-dots interpolant $f_{{\mathcal{D}}}$ is always a solution to (3), since we may take any solution $f$ of (3) and remove all knots from it (resulting in the function $f_{{\mathcal{D}}}$) without increasing its representational cost. If ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are aligned for some $i=2,\dots,N-2$, we can view any interpolant on the interval $[x_{i},x_{i+1}]$ as an instance of Figure 3 with ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and $\bm{c}={\bm{s}}_{i+1}$. By Lemma 3.2, any neural network with a knot at some point $\tilde{x}\in(x_{i},x_{i+1})$ can have the same representational cost as the connect-the-dots solution on this interval, only if ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned.
@@
We can also prove by contradiction that optimal solutions are unique on $[x_{i},x_{i+1}]$ as long as ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Suppose that there is some other optimal interpolant $f$ which is not the connect-the-dots solution $f_{{\mathcal{D}}}$ on an interval $[x_{i},x_{i+1}]$ for which ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ point in different directions. Then apply the lemma repeatedly to remove all knots from $f_{\bm{\theta}}$ which are not located at the data points, except for a single remaining knot at some $\tilde{x}$ between consecutive data points. If this knot occurs after $x_{2}$ (the second data point) or before $x_{N-1}$ (second to last data point), the lemma implies automatically (again taking ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and ${\bm{s}}_{i+1}$) that removing this knot would strictly decrease the network’s representational cost, contradicting optimality of $f$. To conclude the proof, it remains only to show that any optimal interpolant of the dataset must agree with the connect-the-dots interpolant $f_{{\mathcal{D}}}$ before $x_{2}$ and after $x_{N-1}$; the details of this argument appear in Appendix 6.1. As our theorem and corollary quantify, real-world regression datasets (which are typically real-valued and often assumed to incorporate some random noise from an absolutely continuous distribution, e.g. Gaussian) are extremely unlikely to satisfy this special alignment condition; hence, our claim that connect-the-dots interpolation is almost always the unique solution to (3).
@@
We provide numerical examples to illustrate the difference in solutions obtained from single task versus multi-task training and validate our theorem. The first row in Figure 4 shows three randomly initialized univariate neural networks trained to interpolate the five red points with minimum representational cost. While all three of the learned functions have the same representational cost (i.e., all minimize the second-order total variation subject to the interpolation constraint), they each learn different interpolants. This demonstrates that gradient descent does not induce a bias towards a particular solution. The second row shows the function learned for the first output of a multi-task neural network. This network was trained on two tasks. The first task consists of interpolating the five red points while the second consists of interpolating five randomly generated labels sampled from a standard Gaussian distribution. When trained to minimum representation cost we see that the connect-the-dots solution is the only solution learned regardless of initialization, verifying Theorem 3.1. This solution simultaneously minimizes the second-order total variation and (on the interval $[x_{1},x_{N}]$) the norm in the first-order Sobolev RKHS $H^{1}([x_{1},x_{N}])$ associated with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966), subject to the interpolation constraints.
Figure 4: Top Row: Three randomly initialized neural networks trained to interpolate the 5 red points with minimum sum of squared weights. Bottom Row: Three randomly initialized two-output neural networks trained to interpolate a multi-task dataset with minimum sum of squared weights. The labels for the first task are the 5 red points shown while the labels for the second were randomly sampled from a standard Gaussian distribution.

### v3

@@ Abstract
This paper studies the properties of solutions to multi-task shallow ReLU neural network learning problems, wherein the network is trained to fit a dataset with minimal sum of squared weights. Remarkably, the solutions learned for each individual task resemble those obtained by solving a kernel regression problem, revealing a novel connection between neural networks and kernel methods. It is known that single-task neural network learning problems are equivalent to a minimum norm interpolation problem in a non-Hilbertian Banach space, and that the solutions of such problems are generally non-unique. In contrast, we prove that the solutions to univariate-input, multi-task neural network interpolation problems are almost always unique, and coincide with the solution to a minimum-norm interpolation problem in a Sobolev (Reproducing Kernel) Hilbert Space. We also demonstrate a similar phenomenon in the multivariate-input case; specifically, we show that neural network learning problems with large numbers of tasks are approximately equivalent to an $\ell^{2}$ (Hilbert space) minimization problem over a fixed kernel determined by the optimal neurons.
@@ 1 Introduction
This paper characterizes the functions learned by multi-output shallow ReLU neural networks trained with weight decay regularization, wherein each network output fits a different “task” (i.e., a different set of labels on the same data points). We show that the solutions to such multi-task training problems can differ dramatically from those obtained by fitting separate neural networks to each task individually. Unlike standard intuitions Caruana (1997) and existing theory Ben-David and Schuller (2003); Maurer et al. (2016) regarding the effects and benefits of multi-task learning, our results do not rely on similarity between tasks.
@@
Neural networks are trained to fit data using gradient descent methods and often include a form of regularization called weight decay, which penalizes the $\ell^{2}$ norm of the network weights. We consider weight decay applied only to the input and output weights of the neurons—no regularization is applied to the biases or residual connection. This is a common setting studied frequently in past work Savarese et al. (2019); Ongie et al. (2019); Parhi and Nowak (2021). Intuitively, only the input and output weights—not the biases or residual connection—affect the “regularity” of the neural network function as measured by its second (distributional) derivative, which is why it makes sense to regularize only these parameters. Given a set of training data points $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$ and a fixed width $K\geq N^{2}$, we consider the weight decay interpolation problem:  [^1] By an argument similar to the proof of Theorem 5 of Shenouda et al. (2024), as long as $K\geq N^{2}$, problem (3) is well-posed and attains the same minimal objective value (regardless of which $K\geq N^{2}$ is chosen). Therefore, in this work, we always assume that $K$ is some fixed value larger than $N^{2}$.
@@
The optimizations in (2) and (3) are non-convex and in general, they may have multiple global minimizers. As an example, consider the single-task, univariate dataset in Fig. 1. For this dataset, (3) has infinitely many global solutions Savarese et al. (2019); Ergen and Pilanci (2021); Debarre et al. (2022); Hanin (2021). Two of the global minimizers are shown in Fig. 1. In some scenarios, the solution on the right may be preferable to the one on the left, since the interpolation function stays closer to the training data points, and hence is more adversarially robust by most definitions Carlini et al. (2019). Moreover, recent theoretical work shows that this solution has other favorable generalization and robustness properties Joshi et al. (2024). Current training methods, however, might produce any one of the infinite number of solutions, depending on the random initialization of neural network weights as well as other possible sources of randomness in the training process. It is impossible to control this using existing training algorithms, which might explain many problems associated with current neural networks such as their sensitivity to adversarial attacks. In contrast, as we show in this paper, if we train a neural network to interpolate the data in Fig. 1 along with additional interpolation tasks with different labels, then the solution is almost always unique and given by the (potentially preferable) interpolation depicted on the right. This demonstrates that multi-task learning can have a profound effect on the nature of neural network solutions.
@@
  2. Multi-Task Learning $\equiv$ Kernel Method (almost always).
    In the case that the solution to the univariate weight decay problem is unique, it is given by the connect-the-dots interpolant of the training data points: i.e., the optimal solution is a linear spline which performs straight-line interpolation between consecutive data points in all tasks. On the support of the data, this solution agrees with the solution to minimum-norm interpolation in the first-order Sobolev space $H^{1}$, a reproducing kernel Hilbert space (RKHS) which contains all functions with first derivatives in $L^{2}$ De Boor and Lynch (1966). In contrast, solutions to the single-task learning problem are non-unique in general and are given by minimum-norm interpolating functions in the non-Hilbertian Banach space $\operatorname{BV}^{2}$ Parhi and Nowak (2021), which contains all functions with second distributional derivatives [^2] Technically, $\operatorname{BV}^{2}$ contains all functions with second distributional derivatives in ${\mathcal{M}}$, the space of Radon measures with finite total variation. ${\mathcal{M}}$ can be viewed as a “generalization” of $L^{1}$ (see Ch. 7.3, p.223 in Folland (1999)). in $L^{1}$. This shows that the solution to each individual task in a multi-task solution is equivalent to that of a minimum-norm kernel interpolation problem, whereas single-task solutions generally are not.
  3. Insights on Multivariate Multi-Task Learning.
    We provide empirical evidence and mathematical analysis which indicate that similar conclusions hold in multivariate settings. Specifically, with a large number of tasks, the solutions for each individual task in multi-task learning are approximately minimum-norm solutions in a particular RKHS determined by the optimal neurons. In contrast, learning each task in isolation results in solutions that are minimum-norm with respect to a non-Hilbertian Banach norm over the optimal neurons.
@@ 2 Related Works
Characterizations of ReLU neural network solutions:
Hanin (2021); Stewart et al. (2023) characterized the neural network solutions to (2) in the univariate input/output setting. Boursier and Flammarion (2023) showed that in the univariate input/output case, when weight decay is modified to include the biases of each neuron, the solution is unique. Moreover, under certain assumptions, it is often the sparsest interpolant (i.e., the interpolant with the fewest neurons). Our work differs from these in that we study the multi-task setting, showing that univariate multi-task solutions are almost always unique and equivalent to the connect-the-dots solution, which is generally not the sparsest, and is a minimum-norm solution in a Sobolev RKHS. While characterizing solutions to (2) in the multivariate setting is more challenging, there exist some results for very particular datasets Ergen and Pilanci (2021); Ardeshir et al. (2023); Zeno et al. (2024).
@@
For single-output ReLU neural networks, Savarese et al. (2019); Ongie et al. (2019) related weight decay regularization on the parameters of the model to regularizing a particular semi-norm on the neural network function. Ongie et al. (2019) showed that this semi-norm is not an RKHS semi-norm, highlighting a fundamental difference between learning with neural networks and kernel methods. Parhi and Nowak (2021, 2022); Bartolucci et al. (2023); Unser (2021) studied the function spaces associated with this semi-norm, and developed representer theorems showing that optimal solutions to the minimum-norm data fitting problem over these spaces are realized by finite-width ReLU networks. Consequently, finite-width ReLU networks trained with weight decay are optimal solutions to the regularized data-fitting problem posed over these spaces. Function spaces and representer theorems for multi-output and deep neural networks were later developed in Korolev (2022); Parhi and Nowak (2022); Shenouda et al. (2024).
@@
The advantages of multi-task learning has been extensively studied in the machine learning literature Obozinski et al. (2006, 2010); Argyriou et al. (2006, 2008); Caruana (1997). Theoretical properties of multi-task neural networks in particular, have also been studied in Lindsey and Lippl (2023); Collins et al. (2024); Shenouda et al. (2024). The underlying intuition in these past works has been that learning multiple related tasks simultaneously can help select or learn the most useful features for all tasks. Our work differs from this traditional paradigm as we consider multi-task neural networks trained on very general tasks which may be diverse and unrelated.
@@ 3 Univariate Multi-Task Neural Network Solutions
For any function $f$ that can be represented by a neural network (1) with width $K$, we define its representational cost to be
$R(f):=\inf_{\bm{\theta}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}\ \quad\text{subject to}\quad\|{\bm{w}}_{k}\|_{2}=1\ \forall k,\quad f=f_{{\bm{\theta}}}$ (4)
@@
denote the slope of the straight line between $(x_{i},y_{it})$ and $(x_{i+1},y_{i+1t})$. The connect-the-dots interpolant of the dataset ${\mathcal{D}}_{t}$ is the function $f_{{\mathcal{D}}_{t}}$ which connects the consecutive points in dataset ${\mathcal{D}}_{t}$ with straight lines (see Fig. 2). Its slopes on $(-\infty,x_{2}]$ and $[x_{N-1},\infty)$ are $s_{1t}$ and $s_{N-1t}$, respectively.
@@
This result raises an interesting new connection between data fitting with ReLU neural networks and traditional kernel-based learning methods. Indeed, connect-the-dots interpolation is also the minimum-norm interpolant over the first-order Sobolev space $H^{1}([x_{1},x_{N}])$, itself an RKHS whose norm penalizes the $L^{2}$ norm of the derivative of the function. In particular, $f_{{\mathcal{D}}_{t}}$ agrees on $[x_{1},x_{N}]$ with the function $f(x)=\sum_{j=1}^{N}\alpha_{j}k(x,x_{j})$ whose coefficients $\alpha_{j}$ solve the kernel optimization problem
@@ 3.1 Characterization and Uniqueness
Previous works Shenouda et al. (2024) and Lindsey and Lippl (2023) showed that multi-task learning encourages neuron sharing, where all task are encouraged to utilize the same set of neurons or representations. Our result above shows that univariate multi-task training is an extreme example of this phenomenon, since $f_{{\mathcal{D}}}$ can be represented using only $N-1$ neurons, all of which contribute to all of the network outputs. Therefore, in the scenario we study here, neuron sharing almost always occurs even if the tasks are unrelated.
The full proof of Theorem 3.1 appears in Section A.1. We outline the main ideas here. Our proof relies on the fact that any $\mathbb{R}\to\mathbb{R}^{T}$ ReLU neural network of the form (1) which solves (3) represents $T$ continuous piecewise linear (CPWL) functions, where the change in slope of the $t^{\textrm{th}}$ function at the $k^{\text{th}}$ knot is equivalent the $t^{\textrm{th}}$ entry of the $k^{\text{th}}$ output weight vector (see Section A.1 for further detail). This fact allows us to draw a one-to-one correspondence between each knot in the function and each neuron in the neural network. The proof relies primarily on the following lemma:
@@
Let $f:\mathbb{R}\to\mathbb{R}^{T}$ be a function for which each output $f_{t}$ is CPWL and interpolates ${\mathcal{D}}_{t}$. Suppose that at some $\tilde{x}$ between consecutive data points, one or more of the outputs $f_{t}$ has a knot. Let $\tilde{x}_{1}$ and $\tilde{x}_{2}$ be the $x$-coordinates of the closest knots before and after $\tilde{x}$, respectively. Denote the slopes of $f_{t}$ around this interval in terms of $a_{t}$, $b_{t}$, $c_{t}$, and $\delta_{t}$ as in Fig. 3, and let ${\bm{a}},{\bm{b}},\bm{c},\bm{\delta}\in\mathbb{R}^{T}$ be the vectors containing the respective values for each task.
@@
Figure 3: The function output $f_{t}$ around the knot at $\tilde{x}$, where $\tau=\frac{\tilde{x}-\tilde{x}_{1}}{\tilde{x}_{2}-\tilde{x}_{1}}$. Each line segment in the figure is labeled with its slope. For any particular output $t$, it may be the case that $f_{t}$ does not have a knot at $\tilde{x}$ (in which case $\delta_{t}=0$); does not have a knot at $\tilde{x}_{1}$ (in which case $a_{t}=b_{t}+\delta_{t}$); and/or does not have a knot at $\tilde{x}_{2}$ (in which case $b_{t}-\frac{\tau}{1-\tau}\delta_{t}=c_{t}$).
@@
When represented by a ReLU neural network each knot in the function corresponds to a neuron. The representational cost (4) is separable across neurons. The contribution of these knots to $R(f)$ is:
@@
where the inequality follows from the triangle inequality. This shows that taking $\delta_{t}=0$ for all outputs, which corresponds to connecting $\tilde{x}_{1}$ and $\tilde{x}_{2}$ with a straight line in all outputs, will never increase the representational cost of $f$. The triangle inequality used in (3.1) holds with equality for some $\bm{\delta}\neq\bm{0}$ if and only if ${\bm{a}}-{\bm{b}}$, ${\bm{b}}-\bm{c}$, and $\bm{\delta}$ are aligned with $\|\bm{\delta}\|_{2}\leq\min\{\|{\bm{a}}-{\bm{b}}\|_{2},\frac{1-\tau}{\tau}\|{\bm{b}}-\bm{c}\|_{2}\}$. ∎
@@
Lemma 3.2 states that removing neurons which are located away from the data points and replacing them with a straight line will never increase the representational cost of the network, and it will strictly decrease the representational cost unless ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned. This result implies that the connect-the-dots interpolant $f_{{\mathcal{D}}}$ is always a solution to (3), since we may take any solution $f$ of (3) and remove all knots from it (resulting in the function $f_{{\mathcal{D}}}$) without increasing its representational cost. If ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are aligned for some $i=2,\dots,N-2$, we can view any interpolant on the interval $[x_{i},x_{i+1}]$ as an instance of Fig. 3 with ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and $\bm{c}={\bm{s}}_{i+1}$. By Lemma 3.2, any neural network with a knot at some point $\tilde{x}\in(x_{i},x_{i+1})$ can have the same representational cost as the connect-the-dots solution on this interval, only if ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned.
@@
We can also prove by contradiction that optimal solutions are unique on $[x_{i},x_{i+1}]$ as long as ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Suppose that there is some other optimal interpolant $f$ which is not the connect-the-dots solution $f_{{\mathcal{D}}}$ on an interval $[x_{i},x_{i+1}]$ for which ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Then apply the lemma repeatedly to remove all knots from $f_{\bm{\theta}}$ which are not located at the data points, except for a single remaining knot at some $\tilde{x}$ between consecutive data points. If this knot occurs after $x_{2}$ (the second data point) or before $x_{N-1}$ (second to last data point), the lemma implies automatically (again taking ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and ${\bm{s}}_{i+1}$) that removing this knot would strictly decrease the representational cost of network, contradicting optimality of $f$. To conclude the proof, it remains only to show that any optimal interpolant of the dataset must agree with the connect-the-dots interpolant $f_{{\mathcal{D}}}$ before $x_{2}$ and after $x_{N-1}$; the details of this argument appear in Section A.1. As our theorem and corollary quantify, real-world regression datasets (which are typically real-valued and often assumed to incorporate some random noise from an absolutely continuous distribution, e.g. Gaussian) are extremely unlikely to satisfy this special alignment condition; hence, our claim that connect-the-dots interpolation is almost always the unique solution to (3).
@@
We provide numerical examples to illustrate the difference in solutions obtained from single task versus multi-task training and validate our theorem. The first row in Fig. 4 shows three randomly initialized univariate neural networks trained to interpolate the five red points with minimum sum of squared weights. While all three of the learned functions have the same representational cost (i.e., all minimize the second-order total variation subject to the interpolation constraint), they each learn different interpolants. This demonstrates that gradient descent does not induce a bias towards a particular solution. The second row shows the function learned for the first output of a multi-task neural network. This network was trained on two tasks. The first task consists of interpolating the five red points while the second consists of interpolating five randomly generated labels sampled from a standard Gaussian distribution. When trained to interpolate with minimum sum of squared weights we see that the connect-the-dots solution is the only solution learned regardless of initialization, verifying Theorem 3.1. This solution simultaneously minimizes the second-order total variation and the norm in the first-order Sobolev RKHS $H^{1}([x_{1},x_{N}])$ associated with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966), subject to the interpolation constraints. [^4] Code for reproducing all numerical experiments can be found at https://github.com/joeshenouda/effects-mtl-nns.
Figure 4: Top Row: Three randomly initialized neural networks trained to interpolate the five red points with minimum sum of squared weights. Bottom Row: Three randomly initialized two-output neural networks trained to interpolate a multi-task dataset with minimum sum of squared weights. The labels for the first task are the five red points shown while the labels for the second were randomly sampled from a standard Gaussian distribution.